OpenStudy (anonymous):

In a G.P. , the first term is 7, the last term is 448, and the sum is 889. Find the common ratio and the number of terms.

OpenStudy (anonymous):

use an=a1*r^(n-1)

OpenStudy (anonymous):

where an=48 and a1=7

OpenStudy (anonymous):

*448

OpenStudy (anonymous):

it is coming r^(n-1)=64

OpenStudy (anonymous):

then i used the sum formula Sn=\[Sn =\frac{ a[r^n-1) }{ r-1 }\]

OpenStudy (anonymous):

exactly...now you got two equations in two unknowns...so solve for r and n

OpenStudy (anonymous):

i am stuck here

OpenStudy (anonymous):

where??tell me the second one

OpenStudy (anonymous):

\[127=\frac{ r^n-1 }{ r-1 }\]

OpenStudy (anonymous):

r comes out to be 2.isn't it??

OpenStudy (anonymous):

The answer should come r=2 and n=7

OpenStudy (anonymous):

yeah hold on lemme write

OpenStudy (anonymous):

ok

OpenStudy (anonymous):